# Automatic Speech Recognition of African American English: Lexical and Contextual Effects

### v1

@@ 1.1.1 Consonant cluster reduction
[…]
@@ 1.1.2 ING-reduction
[…]
@@ 1.2.1 Lexical neighborhood effect
[…]
@@ 2.2 Feature extraction
To extract CCR and ING-reduction features, we employed forced alignment, an approach inspired by Kendall et al.[6]. Their study compared human coding of the sociolinguistic variable (ING) with force alignment and machine learning classifiers, demonstrating that automated coding algorithms can perform close to human coders in their ability to categorize the ING variation. Following this lead, we utilized the Montreal Forced Aligner (MFA, version 2.2.17) [18] and the Carnegie Mellon University (CMU) Pronouncing Dictionary to automate the feature extraction process for our analysis. We identified missing words in our dataset, trained a grapheme-to-phoneme (g2p) model based on the CMU dictionary, and generated pronunciations for these words. We then updated the CMU dictionary with these additions. For words prone to CCR or ING variation, we included both original and reduced pronunciations in the dictionary (e.g., “accept” was represented as both “AH0 K S EH1 P T” and “AH0 K S EH1 P”). Using MFA’s train command, we developed a custom acoustic model on our entire DCA audio set. Finally, we aligned the complete audio set using this trained acoustic model and the updated CMU dictionary.
@@ 2.3 ASR transcription
We employed wav2vec 2.0 [19] as one of the end-to-end ASR models to transcribe our data, specifically using the pretrained model facebook/wav2vec2-large-960h. This version, trained on 960 hours of speech, was subsequently enhanced with an external 5-gram LM trained on CORAAL’s DCA and DCB subcorpora (entire Washington DC data) using KenLM [20]. To enable transcription with and without LM, we first resampled our audio files to 16 kHz for compatibility with wav2vec 2.0. The audio was then segmented into chunks of at least 30 seconds, focusing on CCR and ING-reduction prone words, while ensuring no sentences were split mid-utterance by leveraging CORAAL’s provided time frames for speaker utterances. The segmented audio chunks were processed through wav2vec 2.0 both with and without the LM. Subsequently, we aligned the transcriptions with their corresponding ground truth using the Needleman-Wunsch algorithm for sequence alignment, implemented via the Python string2string [^2] https://github.com/stanfordnlp/string2string library. This alignment allowed us to evaluate transcription accuracy.
@@ 2.4 Post processing
Following Luce’s [21] definition, we considered a word a lexical neighbor if it could be derived from the target word through a single phoneme substitution, deletion, or addition in any position. To identify these neighbors, we employed the Levenshtein algorithm [22]. Our process involved first checking if the ASR transcribed word existed in the CMU dictionary. For words not found, we generated pronunciations using MFA’s g2p command, and then updated the CMU dictionary with the new entries. We then calculated the phonological Levenshtein distance to compile a list of lexical neighbors for each target word. We used the MFA_Status of each target word to determine whether it was detected by MFA in its original or reduced form. If original, we generated the list of neighbors according to the word’s full pronunciation. Otherwise (for reduced form detection), we obtained the list of neighbors based on the word’s reduced pronunciation. Eventually, if the transcribed word appeared among these neighbors, we attributed the transcription error to the lexical neighborhood effect.
@@ 3.1.2 Statistical procedure
Linear mixed-effects models were fitted using lmer function from the lmerTest package in R (Version 4.4.1) [23]. The models specify that WER is predicted by the fixed effects of MFA_Status, AgeGroup, and Gender, with random intercepts and slopes for these predictors across Target_Word, and random intercepts and slopes for MFA_Status across Speaker_Id. [^3] The model formula: WER $\sim$ MFA_Status + AgeGroup + Gender + (1 + MFA_Status + AgeGroup + Gender | Target_Word) + (1 + MFA_Status | Speaker_Id). The categorical variables were contrast coded. MFA_Status was sum coded as $-0.5$ for “Original Pronunciation” and $0.5$ for “Reduced Pronunciation”. Gender was coded as $-0.5$ for male and $0.5$ for female. AgeGroup was Helmert coded to compare each level with the mean of the previous levels.
@@ 3.2.3 Summary of the results
In ASR without an LM, we observed 7.9% (1,006) of neighborhood errors out of 12,734 total incorrect transcriptions for our target words. However, with the integration of an LM, the number of neighborhood errors drastically decreased to 3.3% (277) out of the total misrecognitions of 8,283. This descriptive finding is confirmed by the regression model, which reveals significant effects across all datasets, indicating that language model usage influences lexical neighborhood errors. ASR_Type shows a consistent, significant positive effect ($p$s $<0.001$) when comparing ASR with and without LM. This effect is strongest for the ING dataset $(\hat{\beta}:-2.1879)$, followed by the overall dataset $(\hat{\beta}:-1.2875)$, and the CCR dataset $(\hat{\beta}:-0.8954)$.
@@ 4 Discussion
In our second hypothesis, we argued that integrating an external LM into the ASR model would reduce errors stemming from lexical neighborhood effect. This was strongly supported by our findings in Section 3.2.3. In other words, while end-to-end ASR models are often promoted for their ability to eliminate the need for separate LMs [30], our results align with recent research [12, 11, 14] that underscores the continued importance of LMs in improving ASR performance. As also illustrated in Figure 1, incorporating an LM significantly reduced WER for both the CCR and ING datasets. This reduction can be attributed to the LM’s ability to provide contextual predictability, thereby mitigating the lexical neighborhood effect.
@@ 5 Conclusion
This study examined the performance of ASR systems, focusing on CCR and ING-reduction, two common phonological variations in AAE. Our findings underscore the persistent challenges ASR systems face when transcribing dialectal speech, even with advanced architectures like wav2vec 2.0 and the integration of LMs. First, our results confirmed that CCR and ING-reduction variations significantly contribute to ASR misrecognitions, strongly supporting our initial hypothesis. To further explore this, we analyzed the effects of gender and age on ASR performance. While gender showed no significant impact, age was a critical factor among AAE speakers under 19, leading to higher rates of ASR misrecognition. Second, across all datasets, ASR with an LM consistently outperformed the one without an LM in reducing neighborhood errors. This validates our second hypothesis and highlights the LM’s ability to leverage contextual predictability, minimize confusion between phonetically similar words, and improve transcription accuracy.
@@
The work is part of HM’s PhD. Conceptualization, Methodology, Formal Analysis, Writing – original draft/review/editing: HM, KT; Data Curation, Investigation, Funding Acquisition, Project Administration, Software, Validation: HM; Resources, Supervision: KT.

### v2

@@ 1.1.1 Consonant cluster reduction
In the error classification conducted by Wassink et al. [3], consonant cluster reduction was identified as the most frequent feature contributing to AAE-related errors. However, the specific origins of these errors remain underexplored. CCR is defined as the simplification of word-final consonant clusters, typically involving the omission of the final stop in a cluster of two consonants (e.g., cold /koUld/ → [koUl]) or the penultimate consonant in a cluster of three (e.g., fists /fIsts/ → [fIs:]) [4]. Given the greater prevalence of two-consonant clusters compared to three-consonant clusters in English [5], this study concentrates specifically on two-consonant clusters ending in a final stop.
@@ 1.1.2 ING-reduction
In sociolinguistic research, ING-reduction refers to pronunciation variation in words ending with -ing, focusing on the realization of the final nasal segment [6]. This variation manifests in two primary forms: the standard -ing pronunciation with a velar nasal [N] and the reduced -in pronunciation with an alveolar nasal [n]. It occurs both within individual morphemes (e.g., the progressive suffix -ing) and within larger word forms (e.g., something, during), while monosyllabic words (e.g., thing, king) are excluded as they do not exhibit this variability.
@@ 1.2.1 Lexical neighborhood effect
Phonological reduction, where words are pronounced in a shortened or simplified form, can lead to non-word percepts that are more prone to misrecognition. These reduced forms often have denser lexical neighborhoods due to their shorter length, which increases the challenge of accurate perception [10]. For instance, the reduced form of ``test'' [tEs] has 21 neighbors, including words like `guess' and `ten', as identified using the CLEARPOND database [10]. This interaction between phonological reduction and lexical neighborhood density underscores the complex challenges faced by both human listeners and ASR systems in accurately perceiving speech, particularly in casual or fast-paced conversational contexts [7, 9].
@@ 2.2 Feature extraction
To extract CCR and ING-reduction features, we employed forced alignment, an approach inspired by Kendall et al.[6]. Their study compared human coding of the sociolinguistic variable (ING) with force alignment and machine learning classifiers, demonstrating that automated coding algorithms can perform close to human coders in their ability to categorize the ING variation. Following this lead, we utilized the Montreal Forced Aligner (MFA, version 2.2.17) [18] and the Carnegie Mellon University (CMU) Pronouncing Dictionary to automate the feature extraction process for our analysis. We identified missing words in our dataset, trained a grapheme-to-phoneme (g2p) model based on the CMU dictionary, and generated pronunciations for these words. We then updated the CMU dictionary with these additions. For words prone to CCR or ING variation, we included both original and reduced pronunciations in the dictionary (e.g., ``accept'' was represented as both ``AH0 K S EH1 P T'' and ``AH0 K S EH1 P''). Using MFA's train command, we developed a custom acoustic model on our entire DCA audio set. Finally, we aligned the complete audio set using this trained acoustic model and the updated CMU dictionary.
@@ 2.3 ASR transcription
We employed wav2vec 2.0 [19] as one of the end-to-end ASR models to transcribe our data, specifically using the pretrained model facebook/wav2vec2-large-960h. This version, trained on 960 hours of speech, was subsequently enhanced with an external 5-gram LM trained on CORAAL's DCA and DCB subcorpora (entire Washington DC data) using KenLM [20]. To enable transcription with and without LM, we first resampled our audio files to 16 kHz for compatibility with wav2vec 2.0. The audio was then segmented into chunks of at least 30 seconds, focusing on CCR and ING-reduction prone words, while ensuring no sentences were split mid-utterance by leveraging CORAAL’s provided time frames for speaker utterances. The segmented audio chunks were processed through wav2vec 2.0 both with and without the LM. Subsequently, we aligned the transcriptions with their corresponding ground truth using the Needleman-Wunsch algorithm for sequence alignment, implemented via the Python string2string [^2] https://github.com/stanfordnlp/string2string library. This alignment allowed us to evaluate transcription accuracy.
@@ 2.4 Post processing
Following Luce's [21] definition, we considered a word a lexical neighbor if it could be derived from the target word through a single phoneme substitution, deletion, or addition in any position. To identify these neighbors, we employed the Levenshtein algorithm [22]. Our process involved first checking if the ASR transcribed word existed in the CMU dictionary. For words not found, we generated pronunciations using MFA's g2p command, and then updated the CMU dictionary with the new entries. We then calculated the phonological Levenshtein distance to compile a list of lexical neighbors for each target word. We used the MFA_Status of each target word to determine whether it was detected by MFA in its original or reduced form. If original, we generated the list of neighbors according to the word's full pronunciation. Otherwise (for reduced form detection), we obtained the list of neighbors based on the word's reduced pronunciation. Eventually, if the transcribed word appeared among these neighbors, we attributed the transcription error to the lexical neighborhood effect.
@@ 3.1.2 Statistical procedure
Linear mixed-effects models were fitted using lmer function from the lmerTest package in R (Version 4.4.1) [23]. The models specify that WER is predicted by the fixed effects of MFA_Status, AgeGroup, and Gender, with random intercepts and slopes for these predictors across Target_Word, and random intercepts and slopes for MFA_Status across Speaker_Id. [^3] The model formula: WER $\sim$ MFA_Status + AgeGroup + Gender + (1 + MFA_Status + AgeGroup + Gender | Target_Word) + (1 + MFA_Status | Speaker_Id). The categorical variables were contrast coded. MFA_Status was sum coded as 0.5-0.5 for ``Original Pronunciation'' and $0.5$ for ``Reduced Pronunciation''. Gender was coded as 0.5-0.5 for male and $0.5$ for female. AgeGroup was Helmert coded to compare each level with the mean of the previous levels.
@@ 3.2.3 Summary of the results
In ASR without an LM, we observed 7.9% (1,006) of neighborhood errors out of 12,734 total incorrect transcriptions for our target words. However, with the integration of an LM, the number of neighborhood errors drastically decreased to 3.3% (277) out of the total misrecognitions of 8,283. This descriptive finding is confirmed by the regression model, which reveals significant effects across all datasets, indicating that language model usage influences lexical neighborhood errors. ASR_Type shows a consistent, significant positive effect ($p$s <0.001<0.001) when comparing ASR with and without LM. This effect is strongest for the ING dataset $(\hat{\beta}:-2.1879)$, followed by the overall dataset $(\hat{\beta}:-1.2875)$, and the CCR dataset $(\hat{\beta}:-0.8954)$.
@@ 4 Discussion
In our second hypothesis, we argued that integrating an external LM into the ASR model would reduce errors stemming from lexical neighborhood effect. This was strongly supported by our findings in Section 3.2.3. In other words, while end-to-end ASR models are often promoted for their ability to eliminate the need for separate LMs [30], our results align with recent research [12, 11, 14] that underscores the continued importance of LMs in improving ASR performance. As also illustrated in Figure 1, incorporating an LM significantly reduced WER for both the CCR and ING datasets. This reduction can be attributed to the LM's ability to provide contextual predictability, thereby mitigating the lexical neighborhood effect.
@@ 5 Conclusion
This study examined the performance of ASR systems, focusing on CCR and ING-reduction, two common phonological variations in AAE. Our findings underscore the persistent challenges ASR systems face when transcribing dialectal speech, even with advanced architectures like wav2vec 2.0 and the integration of LMs. First, our results confirmed that CCR and ING-reduction variations significantly contribute to ASR misrecognitions, strongly supporting our initial hypothesis. To further explore this, we analyzed the effects of gender and age on ASR performance. While gender showed no significant impact, age was a critical factor among AAE speakers under 19, leading to higher rates of ASR misrecognition. Second, across all datasets, ASR with an LM consistently outperformed the one without an LM in reducing neighborhood errors. This validates our second hypothesis and highlights the LM's ability to leverage contextual predictability, minimize confusion between phonetically similar words, and improve transcription accuracy.
@@
The work is part of HM's PhD. Conceptualization, Methodology, Formal Analysis, Writing – original draft/review/editing: HM, KT; Data Curation, Investigation, Funding Acquisition, Project Administration, Software, Validation: HM; Resources, Supervision: KT.